substitution method
y=13-3x
2x+9y=-8

Answer by fazlerabbi(9)

Given,
y = 13 - 3x ----- (1)
2x + 9y = -8 ----- (2)
Substituting value of y in (2) from (1) we get
2x+9(13-3x) = -8
=> 2x+117-27x = -8
=> 117-25x = -8
=> 117-25x-117 = -8-117 ;(adding -117 to both sides)
=> -25x = -125
=> x = 5 ;(multiplying both sides by 1/-25)
Substituting the value of x in (1) we get
y = 13 - 3(5)
=>y = 13 - 15
=>y = -2
So, Answer is x=5, y=-2